In the system shown, a horizontal force Fx = 40 N acts on an object of mass m2 = 8.00 kg. A friction force fk = 5 N acts between object m2 and the table top, and m1 = 2.00 kg. i) Calculate the acceleration of the masses. ii) Calculate the tension in the string.
1
Expert's answer
2020-06-01T14:18:43-0400

Given,

The masses "m_1=2Kg,m_2=8Kg" and frictional force acting between table and those masses is "f=f_k=5N" .

Horizontal external force is "F_x=40N" acting on object of mass "m_2" and also both masses are connected by string.

Now, lets draw the Free Body Diagram(FBD) for the system.


i). Now applying Newton's second law we get

"F_{net}=m_{total}\cdot a_{net}\\
\implies (F_x-f_k)=(m_1+m_2)a_{net}\\
\implies 40-5=(2+8)a_{net}\implies a_{net}=3.5m/s^2"

Thus, acceleration of each mass is "3.5m\/s^2" .


ii). Let "T" denotes the tension force.

Now,consider the FBD object of mass "m_1"


Clearly, from the above diagram we get,

"T-f_k=m_1a_{net}\\
\implies T=f_k+m_1a_{net}\implies T=5+2\times3.5=12N"

Thus, Tension on the string is "12N" .